Diving boards and platforms offer a nearly ideal opportunity in which to experience the various laws of motion. When you jump off the high diving board, you are a falling object and, if you can keep your presence of mind as you fall, you will be demonstrating laws of physics. Imagine yourself diving off a platform 10 m above the water below.

How high would the board need to be for a 175 pound man to reach terminal velocity by the time he hit the water? Why would this be a problem?
Answer #1

The velocity of a man when reaching water will be sqrt(2gh).

So the height should be h = v2/2g = 1045.272/2x9.8 = 55.74 km